# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 6.4. How are LLMs used in software development?
API entity and relation extraction. Extracting APIs and their semantic relationships from unstructured text (e.g., data from Stack Overflow) is a fundamental task in SE, but existing methods require labor-intensive manual rule creation or data labeling. Huang et al. (Huang et al., 2023b) present an innovative approach, AERJE, that leverages LLMs for this task. AERJE consists of a BERT-based dynamic hint generator and a T5-based joint entity-relationship extractor, which together enable efficient extraction of API entities and relationships without manual effort. The approach achieved an F1 score of 96.51% for API entity extraction and 81.2% for API relationship extraction, offering a significant advancement over traditional methods.
Code optimization. Efficiency in programming is vital, particularly in resource-limited or large-scale applications. Traditional optimizing compilers enhance efficiency through various considerations like algorithm and data structure selection (Aho et al., 2007). Madaan et al.(Madaan et al., 2023) explore the use of LLMs in suggesting performance-enhancing code edits. They curate a dataset of Performance-Improving Edits (PIE), showing how Codex and CodeGen can generate these edits, resulting in over 2.5x speedups for more than 25% of the C++ and Python programs, even after C++ code was compiled using the O3 optimization level.
@@
Code example recommendation. Zhou et al. (Zhou et al., 2019) pointed out that software developers tend to write similar code examples several times due to the need to implement similar features in different projects. Therefore, during the software development process, recommender systems can provide programmers with the most pertinent and high-quality examples written by other programmers, thus helping them to complete their tasks quickly and efficiently (Di Rocco et al., 2021). Open-source projects and informal documentation are the two main sources of information that developers rely on to perform programming tasks. For example, open-source projects on GitHub provide code examples and code resources for various tasks. Rahmani et al. (Rahmani et al., 2023) introduce a methodology to improve code example recommendations for Java programming language on Stack Overflow using BERT and Query-Aware Locality-Sensitive Hashing (LSH). They employ BERT to convert code into numerical vectors and then apply two LSH variants, Random Hyperplane-based, and Query-Aware, to identify Approximate Nearest Neighbors (ANN).
@@
Control flow graph generation. Control Flow Graphs (CFGs) are a cornerstone of SE that illustrate program behavior by showing sequences of statements and their execution order conditions  (Allen, 1970). As a graphical representation of program behavior, CFGs are critical in many SE tasks, including code search (Guo et al., 2020; Chen et al., 2019b), code clone detection (Wang et al., 2020a; Hu et al., 2018; Wei and Li, 2017) and code classification (Wang et al., 2020b; Zhang et al., 2019). Huang et al. (Huang et al., 2023d) presented a novel approach for generating behaviorally correct CFGs of statically typed partial code by leveraging the error-tolerant and understanding ability of LLMs. The approach involves a Chain of Thoughts (CoT) with four steps: structure hierarchy extraction, nested code block extraction, CFG generation of nested code blocks, and fusion of all nested code blocks’ CFGs (Le-Cong et al., 2022). The CoT is broken down into an AI chain according to the single responsibility principle, along with effective prompt instructions. This results in superior node and edge coverage compared to traditional program analysis-based methods and the original CoT method.
@@ 6.5. How are LLMs used in software quality assurance?
Within the domain of software quality assurance, LLMs have emerged as valuable tools with diverse applications for various tasks, including vulnerability detection, test generation, bug localization, etc.
Test generation. Test generation involves automating the process of creating test cases to evaluate the correctness and functionality of software applications. It encompasses various aspects, including test case generation (Zhang et al., 2023g), unit test generation (Tang et al., 2023a; Yuan et al., 2023b; Schäfer et al., 2023; Xie et al., 2023; Siddiq et al., 2023b), etc. LLM application in test generation offers several advantages, including the ability to automatically generate diverse test cases, improving test coverage (Schäfer et al., 2023; Siddiq et al., 2023b) and identifying potential defects (Xie et al., 2023). LLMs can also assist in generating test cases based on natural language descriptions, fostering better collaboration between developers and testers. Additionally, they help identify areas lacking test coverage and suggest relevant test cases, ensuring comprehensive testing and reducing the risk of undiscovered issues (Zhang et al., 2023g). By enhancing test efficiency and effectiveness, LLMs contribute to producing more reliable and high-quality software products.
Vulnerability detection. The number of software vulnerabilities is rapidly increasing, as shown by the vulnerability reports from Common Vulnerabilities and Exposures (CVEs) (Anon, 2022) in recent years. As the number of vulnerabilities increases, there will be more possibilities for cybersecurity attacks, which can cause serious economic and social harm. Therefore, vulnerability detection is crucial to ensure the security of software systems and protect social and economic stability. Traditional static detection methods are based on static analysis and predefined matching rules, which rely on developers’ expertise and make it difficult to detect unknown vulnerabilities. With the assistance of LLMs (Thapa et al., 2022; Chan et al., 2023; Chen et al., 2023a), Alqarni et al. (Alqarni and Azim, 2022) present an updated BERT model fine-tuned for vulnerability detection. Additionally, Tang et al. (Tang et al., 2023b) introduced novel approaches using LLMs to enhance vulnerability detection. One of their proposed models, CSGVD, combines sequence and graph embedding for function-level vulnerability detection, outperforming other deep learning-based models on a real-world benchmark dataset. Their study also explores the application of CodeT5 for vulnerability detection, highlighting the importance of code-specific pre-training tasks.
@@
Failure-inducing test identification. Test suites typically include two types of test cases: pass-through test cases and fault-inducing test cases (Li et al., 2023f). In practice, there are far more pass test cases for faults than fault-inducing test cases, which hinders the effectiveness of program debugging. However, in practice, it is difficult to find fault-inducing test cases. This is because developers first need to find test inputs that trigger program faults, and the search space for such test inputs is huge  (Fraser et al., 2015). Moreover, developers need to build a test oracle to automatically detect program faults, and building a test oracle is often an undecidable problem (Ibrahimzada et al., 2022). Li et al. (Li et al., 2023f) investigated the application of ChatGPT to the task of finding fault-inducing test cases in SE. While recognizing ChatGPT’s potential, they initially observed suboptimal performance in pinpointing these cases, particularly when two versions of a program had similar syntax. The authors identified this as a weakness in ChatGPT’s ability to discern subtle code differences. To enhance its performance, they devised a novel approach blending ChatGPT with difference testing. Leveraging ChatGPT’s strength in inferring expected behavior from erroneous programs, they synthesized programs that amplified subtle code differences. The experimental results reveal that this approach greatly increases the probability of finding the correct fault-inducing test case.
Flaky test prediction. In many environments, it has been found that test cases can be non-deterministic, with test cases passing and failing in different executions, even for the same version of the source code. These test cases are called piecewise test cases (Zolfaghari et al., 2021; Luo et al., 2014; Eck et al., 2019; Fatima et al., 2022). Fatima et al. (Fatima et al., 2022) propose a black-box approach named Flakify that uses CodeBERT to predict flaky tests. The model is trained on a dataset of test cases labeled as flaky or non-flaky. The model’s predictions can help developers focus their debugging efforts on a subset of test cases that are most likely to be flaky, thereby reducing the cost of debugging in terms of both human effort and execution time.
@@ 6.6. How are LLMs used in software maintenance?
Program repair. The goal of automated program repair (APR) is to automatically identify and fix bugs or defects in software (Zhang et al., 2023c). It involves leveraging automated techniques to analyze buggy code and generate correct patches to address the identified issues. LLMs, such as BERT (Zhang et al., 2023a; Tian et al., 2023a), CodeBERT (Le-Cong et al., 2023), CodeT5 (Paul et al., 2023a), Codex (Fan et al., 2022; Jin et al., 2023; Wu et al., 2023a), PLBART (Paul et al., 2023a; Wu et al., 2023a), T5 (Yuan et al., 2022; Mastropaolo et al., 2022b) and GPT series (Xia and Zhang, 2023b; Tian et al., 2023b; Xia and Zhang, 2023a; Lajkó et al., 2022; Charalambous et al., 2023; Sobania et al., 2023; Cao et al., 2023), have shown effectiveness in generating syntactically correct and contextually relevant code. Leveraging LLMs for program repair can achieve competitive performance in generating patches for various types of bugs and defects (Xia and Zhang, 2023b). These models can effectively capture the underlying semantics and dependencies in the code (Charalambous et al., 2023), leading to the production of accurate and effective patches (Zhang et al., 2023a; Xia and Zhang, 2023a). Moreover, LLMs can be fine-tuned on specific code repair datasets (Mastropaolo et al., 2022b), further improving their ability to generate high-quality patches for real-world software projects. The application of LLMs in program repair not only accelerates the bug-fixing process but also enables software developers to focus on more complex tasks, leading to enhanced software reliability and maintainability.
@@
Codex GPT-Neo, GPT-J, GPT-NeoX, CodeT5, InCoder QuixBugs-Python and Java, Defects4J 1.2 and 2.0, ManyBugs Correct / plausible patches May 20, 2023 (Xia et al., 2023)
Codex CodeT5, CodeGen, PLBART, InCoder Vul4J, VJBench, Correct / plausible patches May 29, 2023 (Wu et al., 2023a)
ChatGPT Codex, CodeGen-16B, CodeGen-6B, CodeGen-2B, CodeGen-350M QuixBugs-Python and Java Correct / plausible patches Jan 30, 2023 (Xia and Zhang, 2023a)
ChatGPT Codex, CodeBERT, SelfAPR, RewardRepair, Recoder, TBar, CURE, CoCoNuT QuixBugs-Python and Java, Defects4J 1.2 and 2.0 Correct fixes Apr 1, 2023 (Xia and Zhang, 2023b)
@@
  • Domain-specific integration. Merging the capabilities of LLMs with domain-specific knowledge and techniques further enhances their performance. Customized prompts, project-specific fine-tuning, and leveraging SE techniques (Xia et al., 2023; Wang et al., 2023a) can dramatically elevate the efficacy of LLM-driven program repairs.
@@
Debugging. Debugging targets identifying, locating, and resolving software defects or errors, commonly known as bugs. The debugging process involves scrutinizing the code, tracing the execution flow, and isolating the root cause of the problem to effectively correct the error. LLMs, such as BERT and other converter-based architectures, excel at utilizing contextual information and natural language understanding. In terms of debugging, LLMs can be used to simulate the scientific debugging process, such as AutoSD proposed by Kang et al. (Kang et al., 2023). This model generates hypotheses about code problems and extracts relevant values to identify potential problems. In addition, the SELF-DEBUGGING method proposed by Chen et al. (Chen et al., 2023b) enables LLM to debug its own generated code by learning a small number of presentations and explanations, which effectively improves the accuracy and sampling efficiency of code generation. Using LLMs in debugging not only improves fixing performance by generating competitive fixes but also provides insights into and explanations of the model’s decision-making process, making it an important tool for improving software quality and developer productivity.
Bug report analysis. LLMs such as Codex (Kang et al., 2022) and BERT (Ciborowska and Damevski, 2022) comprehensively analyze natural language text, code snippets, and contextual information within bug reports to generate precise code repair suggestions, test cases, or steps for reproducing errors. By deeply understanding the semantics and context of the issues, LLMs offer developers more intelligent solutions, expediting the error-fixing process and alleviating development burdens (Lee et al., 2022; Gomes et al., 2023). These models excel not only in code generation but also in identifying and interpreting crucial information within error reports, aiding developers in better comprehending the underlying causes (Li et al., 2022g). With the integration of LLMs, bug report analysis tasks are conducted more efficiently and accurately advancing optimization and enhancement of the maintenance workflow.
@@
Bug report replay. Bug reports are crucial for software maintenance, allowing users to inform developers of problems encountered while using the software. Therefore, researchers have invested significant resources in automating error playback to speed up the software maintenance process. The success of current automated approaches depends heavily on the characteristics and quality of error reports, as they are limited by manually created schemas and predefined vocabularies. Inspired by the success of the LLMs in natural language understanding, Feng et al. (Feng and Chen, 2023) propose AdbGPT, which utilizes natural language understanding and logical reasoning capabilities of the LLM to extract Steps to Reproduce (S2R) entities from bug reports and guide the bug replay process based on the current graphical user interface (GUI) state. The researchers describe how cue engineering, a small amount of learning, and thought chain reasoning can be utilized to leverage the knowledge of the LLM for automated error replay. This approach is significantly lightweight compared to traditional approaches, which utilize a single LLM to address both phases of S2R entity extraction and guided replay through novel hint engineering.
Duplicate bug report detection. In large software projects, multiple users may encounter and report the same or similar bugs independently, resulting in a proliferation of duplicate bug reports (Isotani et al., 2021). Duplicate bug report detection involves analyzing the textual content of bug reports and comparing them to find similarities and redundancies. LLM models, such as BERT (Isotani et al., 2021), ChatGPT (Sridhara et al., 2023), and other transformer-based architectures, are well-suited for natural language understanding and contextual representation. When applied to this task, LLMs can effectively capture the semantic similarities between bug reports, even in cases with slight variations in language or phrasing. The utilization of LLMs in this context not only enhances efficiency in managing bug reports but also contributes to improving the overall software development and maintenance workflow, reducing redundancy, and ensuring prompt bug resolution (Zhang et al., 2023b).
Decompilation. Decompilation is crucial in many security and SE tasks. For example, decompilation is often the first step in malware analysis (Nafisi, 2021), where human analysts examine malware code to understand its behavior. It is also important for binary vulnerability analysis (where analysts want to identify critical vulnerabilities in executables) (Dinesh et al., 2020; Nagy et al., 2021), software supply chain analysis (Hemel et al., 2011; Ombredanne, 2020), and code reuse (where legacy executables may need to be ported or hardened) (Ding et al., 2019; Marcelli et al., 2022; Pei et al., 2020). Decompilation tools, such as IDA and Ghidra, have been useful in security threat analysis (Nafisi and Lelli, 2021) proves its importance. Xu et al. (Xu et al., 2023) propose a new technique for recovering symbolic names during decompilation that leverages the synergy between LLMs (especially ChatGPT) and program analysis. The method employs an iterative algorithm to propagate ChatGPT query results based on program semantics. This propagation in turn provides better context for ChatGPT. The results show that 75% of the recovered names are perceived as good by the users and that the technique outperforms the state-of-the-art by 16.5% and 20.23% in terms of precision and recall, respectively.

### v6

@@ 6.4. How are LLMs used in software development?
API entity and relation extraction. Extracting APIs and their semantic relationships from unstructured text (e.g., data from Stack Overflow) is a fundamental task in SE, but existing methods require labor-intensive manual rule creation or data labeling. Huang et al. (Huang et al., 2023d) present an innovative approach, AERJE, that leverages LLMs for this task. AERJE consists of a BERT-based dynamic hint generator and a T5-based joint entity-relationship extractor, which together enable efficient extraction of API entities and relationships without manual effort. The approach achieved an F1 score of 96.51% for API entity extraction and 81.2% for API relationship extraction, offering a significant advancement over traditional methods.
@@
Code recommendation. Zhou et al. (Zhou et al., 2019) pointed out that software developers tend to write similar code examples several times due to the need to implement similar features in different projects. Therefore, during the software development process, recommender systems can provide programmers with the most pertinent and high-quality examples written by other programmers, thus helping them to complete their tasks quickly and efficiently (Di Rocco et al., 2021). Open-source projects and informal documentation are the two main sources of information that developers rely on to perform programming tasks. For example, open-source projects on GitHub provide code examples and code resources for various tasks. Rahmani et al. (Rahmani et al., 2023) introduce a methodology to improve code example recommendations for Java programming language on Stack Overflow using BERT and Query-Aware Locality-Sensitive Hashing (LSH). They employ BERT to convert code into numerical vectors and then apply two LSH variants, Random Hyperplane-based, and Query-Aware, to identify Approximate Nearest Neighbors (ANN).
@@
Control flow graph generation. Control Flow Graphs (CFGs) are a cornerstone of SE that illustrate program behavior by showing sequences of statements and their execution order conditions (Allen, 1970). As a graphical representation of program behavior, CFGs are critical in many SE tasks, including code search (Guo et al., 2020; Chen et al., 2019b), code clone detection (Wang et al., 2020a; Hu et al., 2018; Wei and Li, 2017) and code classification (Wang et al., 2020b; Zhang et al., 2019). Huang et al. (Huang et al., 2023g) presented a novel approach for generating behaviorally correct CFGs of statically typed partial code by leveraging the error-tolerant and understanding ability of LLMs. The approach involves a Chain of Thoughts (CoT) with four steps: structure hierarchy extraction, nested code block extraction, CFG generation of nested code blocks, and fusion of all nested code blocks’ CFGs (Le-Cong et al., 2022). The CoT is broken down into an AI chain according to the single responsibility principle, along with effective prompt instructions. This results in superior node and edge coverage compared to traditional program analysis-based methods and the original CoT method.
@@
Others. In addition to the 18 software development tasks detailed above, LLMs can also be applied to code translation (Jana et al., 2023; Pan et al., 2023c, a; Qi et al., 2023; Yan et al., 2023b; Yang et al., 2023g), code editing (Bairi et al., 2023; Gupta et al., 2023; Li et al., 2023e; Moon et al., 2023; Shypula et al., 2023), API documentation augment (Yang et al., 2023c), data analysis (Cheng et al., 2023), fuzz driver generation (Zhang et al., 2023a), instruction generation (Zhou et al., 2023b).
@@ 6.5. How are LLMs used in software quality assurance?
Within the domain of software quality assurance, LLMs have emerged as valuable tools with diverse applications for various tasks, including vulnerability detection, test generation, bug localization, verification, test automation, etc.
Vulnerability detection. The number of software vulnerabilities is rapidly increasing, as shown by the vulnerability reports from Common Vulnerabilities and Exposures (CVEs) (Anon, 2022) in recent years. As the number of vulnerabilities increases, there will be more possibilities for cybersecurity attacks, which can cause serious economic and social harm. Therefore, vulnerability detection is crucial to ensure the security of software systems and protect social and economic stability. Traditional static detection methods are based on static analysis and predefined matching rules, which rely on developers’ expertise and make it difficult to detect unknown vulnerabilities. With the assistance of LLMs (Thapa et al., 2022; Chan et al., 2023; Chen et al., 2023a), Tang et al. (Tang et al., 2023d) introduced novel approaches using LLMs to enhance vulnerability detection. One of their proposed models, CSGVD, combines sequence and graph embedding for function-level vulnerability detection, outperforming other deep learning-based models on a real-world benchmark dataset. Their study also explores the application of CodeT5 for vulnerability detection, highlighting the importance of code-specific pre-training tasks.
Test generation. Test generation involves automating the process of creating test cases to evaluate the correctness and functionality of software applications. It encompasses various aspects, including test case generation (Zhang et al., 2023o), unit test generation (Tang et al., 2023c; Yuan et al., 2023b; Schäfer et al., 2023a; Xie et al., 2023a; Siddiq et al., 2023b), etc. LLM application in test generation offers several advantages, including the ability to automatically generate diverse test cases, improving test coverage (Schäfer et al., 2023a; Siddiq et al., 2023b) and identifying potential defects (Xie et al., 2023a). LLMs can also assist in generating test cases based on natural language descriptions, fostering better collaboration between developers and testers. Additionally, they help identify areas lacking test coverage and suggest relevant test cases, ensuring comprehensive testing and reducing the risk of undiscovered issues (Zhang et al., 2023o). By enhancing test efficiency and effectiveness, LLMs contribute to producing more reliable and high-quality software products.
@@
Fault localization. Test suites typically include two types of test cases: pass-through test cases and fault-inducing test cases (Li et al., 2023l). In practice, there are far more pass test cases for faults than fault-inducing test cases, which hinders the effectiveness of program debugging. However, in practice, it is difficult to find fault-inducing test cases. This is because developers first need to find test inputs that trigger program faults, and the search space for such test inputs is huge (Fraser et al., 2015). Moreover, developers need to build a test oracle to automatically detect program faults, and building a test oracle is often an undecidable problem (Ibrahimzada et al., 2022). Li et al. (Li et al., 2023l) investigated the application of ChatGPT to the task of finding fault-inducing test cases in SE. While recognizing ChatGPT’s potential, they initially observed suboptimal performance in pinpointing these cases, particularly when two versions of a program had similar syntax. The authors identified this as a weakness in ChatGPT’s ability to discern subtle code differences. To enhance its performance, they devised a novel approach blending ChatGPT with difference testing. Leveraging ChatGPT’s strength in inferring expected behavior from erroneous programs, they synthesized programs that amplified subtle code differences. The experimental results reveal that this approach greatly increases the probability of finding the correct fault-inducing test case.
Others. In addition to the six software quality assurance tasks detailed above, LLMs can also be applied to defect detection (Sun et al., 2023a; Wong et al., 2023), GUI testing (Yoon et al., 2023; Liu et al., 2023b), static analysis (Hao et al., 2023; Mohajer et al., 2023), binary taint analysis (Liu et al., 2023f), compiler fuzzing (Quan et al., 2023), decompilation (Xu et al., 2023b), invariant prediction (Pei et al., 2023), malicious code localization (Sun et al., 2023a), mobile app crash detection (Liu et al., 2023c), and resource leak detection (Wang et al., 2023g).
@@ 6.6. How are LLMs used in software maintenance?
Program repair. The goal of automated program repair (APR) is to automatically identify and fix bugs or defects in software (Zhang et al., 2023j). It involves leveraging automated techniques to analyze buggy code and generate correct patches to address the identified issues. LLMs, such as BERT (Zhang et al., 2023c; Tian et al., 2023a), CodeBERT (Le-Cong et al., 2023), CodeT5 (Paul et al., 2023a), Codex (Fan et al., 2022; Jin et al., 2023b; Wu et al., 2023a), PLBART (Paul et al., 2023a; Wu et al., 2023a), T5 (Yuan et al., 2022; Mastropaolo et al., 2022b) and GPT series (Xia and Zhang, 2023b; Tian et al., 2023b; Xia and Zhang, 2023a; Lajkó et al., 2022; Charalambous et al., 2023; Sobania et al., 2023; Cao et al., 2023), have shown effectiveness in generating syntactically correct and contextually relevant code. Leveraging LLMs for program repair can achieve competitive performance in generating patches for various types of bugs and defects (Xia and Zhang, 2023b). These models can effectively capture the underlying semantics and dependencies in the code (Charalambous et al., 2023), leading to the production of accurate and effective patches (Zhang et al., 2023c; Xia and Zhang, 2023a). Moreover, LLMs can be fine-tuned on specific code repair datasets (Mastropaolo et al., 2022b), further improving their ability to generate high-quality patches for real-world software projects. The application of LLMs in program repair not only accelerates the bug-fixing process but also enables software developers to focus on more complex tasks, leading to enhanced software reliability and maintainability.
@@
Codex
GPT-Neo, GPT-J, GPT-NeoX, CodeT5, InCoder
QuixBugs-Python and Java, Defects4J 1.2 and 2.0, ManyBugs
Correct / plausible patches
May 20, 2023 (Xia et al., 2023)
Codex
CodeT5, CodeGen, PLBART, InCoder
Vul4J, VJBench,
Correct / plausible patches
May 29, 2023 (Wu et al., 2023a)
ChatGPT
Codex, CodeGen-16B, CodeGen-6B, CodeGen-2B, CodeGen-350M
QuixBugs-Python and Java
Correct / plausible patches
Jan 30, 2023 (Xia and Zhang, 2023a)
ChatGPT
Codex, CodeBERT, SelfAPR, RewardRepair, Recoder, TBar, CURE, CoCoNuT
QuixBugs-Python and Java, Defects4J 1.2 and 2.0
Correct fixes
Apr 1, 2023 (Xia and Zhang, 2023b)
@@
  • Domain-specific integration. Merging the capabilities of LLMs with domain-specific knowledge and techniques further enhances their performance. Customized prompts, project-specific fine-tuning, and leveraging SE techniques (Xia et al., 2023; Wang et al., 2023c) can dramatically elevate the efficacy of LLM-driven program repairs.
@@
Debugging. Debugging targets identifying, locating, and resolving software defects or errors, commonly known as bugs. The debugging process involves scrutinizing the code, tracing the execution flow, and isolating the root cause of the problem to correct the error effectively. LLMs, such as BERT and other converter-based architectures, excel at utilizing contextual information and natural language understanding. In terms of debugging, LLMs can be used to simulate the scientific debugging process, such as AutoSD proposed by Kang et al. (Kang et al., 2023b). This model generates hypotheses about code problems and extracts relevant values to identify potential problems. In addition, the SELF-DEBUGGING method proposed by Chen et al. (Chen et al., 2023b) enables LLM to debug its own generated code by learning a small number of presentations and explanations, which effectively improves the accuracy and sampling efficiency of code generation. Using LLMs in debugging not only improves fixing performance by generating competitive fixes but also provides insights into and explanations of the model’s decision-making process, making it an important tool for improving software quality and developer productivity.
Bug reproduction. Bug reports are crucial for software maintenance, allowing users to inform developers of problems encountered while using the software. Therefore, researchers have invested significant resources in automating error playback to speed up the software maintenance process. The success of current automated approaches depends heavily on the characteristics and quality of error reports, as they are limited by manually created schemas and predefined vocabularies. Inspired by the success of the LLMs in natural language understanding, Feng et al. (Feng and Chen, 2023) propose AdbGPT, which utilizes natural language understanding and logical reasoning capabilities of the LLM to extract Steps to Reproduce (S2R) entities from bug reports and guide the bug replay process based on the current graphical user interface (GUI) state. The researchers describe how cue engineering, a small amount of learning, and thought chain reasoning can be utilized to leverage the knowledge of the LLM for automated error replay. This approach is significantly lightweight compared to traditional approaches, which utilize a single LLM to address both phases of S2R entity extraction and guided replay through novel hint engineering.
Duplicate bug report detection. In large software projects, multiple users may encounter and report the same or similar bugs independently, resulting in a proliferation of duplicate bug reports (Isotani et al., 2021). Duplicate bug report detection involves analyzing the textual content of bug reports and comparing them to find similarities and redundancies. LLM models, such as BERT (Isotani et al., 2021), ChatGPT (Sridhara et al., 2023), and other transformer-based architectures, are well-suited for natural language understanding and contextual representation. When applied to this task, LLMs can effectively capture the semantic similarities between bug reports, even in cases with slight variations in language or phrasing. The utilization of LLMs in this context not only enhances efficiency in managing bug reports but also contributes to improving the overall software development and maintenance workflow, reducing redundancy, and ensuring prompt bug resolution (Zhang et al., 2023e).
@@
Others. In addition to the 14 software maintenance tasks detailed above, LLMs can also be applied to review/commit/code classification (Ghadhab et al., 2021; Kou et al., 2023a; Yang et al., 2022c), log parsing (Liu et al., 2024b; Ma et al., 2024b; Yu et al., 2023b), code revision (Kabir et al., 2023; Wadhwa et al., 2023), API misuses repair (Zhang et al., 2023h), Code coverage prediction (Tufano et al., 2023), code review explained (Widyasari et al., 2023), Code-Review defects repair (Zhao et al., 2023c), crash bug repair (Du et al., 2023a), dockerfile Repair (Henkel et al., 2021), incivility detection (Ferreira et al., 2024), patch correctness prediction (Zhang et al., 2024a), patch detection (Tang et al., 2023a), rename Refactoring (Liu et al., 2023i), technical debt payback (Mastropaolo et al., 2023a), web test repair (Xu et al., 2023a), type error repair (Chow et al., 2024), etc.